OpenStudy (anonymous):

find the remainder when x51+51 is divided by (x+1).

OpenStudy (anonymous):

is that: \[x^{51}+51\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(which btw, is ridiculous, whoever asked you to do this needs to be punched in the face.)

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

because its to the 51st power. Thats waaaaaay too high for any normal type of problem. \[x^4 +3\] is normal. a polynomial of degree 51 is retarded >.< anywhos, that just my personal feelings on the subject, now for the solution lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you know about synthetic division, or does that term sound familiar to you?

OpenStudy (anonymous):

srry.i dont noe dat

OpenStudy (anonymous):

alright, thats fine, theres is still one more method, i will have to write it out on paper and post it though. One sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if you dont mind me asking, what lvl class is this? algebra II? Calculus?

OpenStudy (anonymous):

nvm, its not that rough of a problem. Basically you do this: \[x^51+51 = (x+1)p(x)+r\] where p(x) is the quotient, and r is the remainder. The (x + 1) is there because we are dividing by (x + 1). If we plug in the value -1 for x we get: \[(-1)^{51}+51 = (-1+1)p(x)+r \Rightarrow -1+51 = r \Rightarrow r = 50\] so the remainder is 50.

OpenStudy (anonymous):

oops, that should be a \[x^{51}\] in the first equation line.